Solve for m. 2/3(m+4)-8=11/3
Answer by checkley77(12844)

2/3(m+4)-8=11/3
2(M+4)/3=11/3+8
(2M+8)/3=11/3+24/3
(2M+8)/3=35/3 CANCEL OUT THE DENOMINATORS.
2M+8=35
2M=35-8
2M=27
M=27/2
M=13.5 ANS.
PROOF:
2/3(13.5+4)-8=11/3
2/3*17.5-8=11/3
35/3-8=11/3
35/3-24/3=11/3
11/3=11/3